40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu May 16 04:46:37 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
    Reply  Thu May 16 23:48:39 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
       Reply  Mon May 20 17:49:36 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data PRMIsensMatQuickAnalysis.m.gz
          Reply  Mon May 20 18:47:47 2013, Koji, Update, LSC, PRMI sensing matrix - not high quality data 
             Reply  Tue May 21 14:48:08 2013, Jenne, Update, LSC, PRMI sensing matrix - not high quality data 
       Reply  Mon May 20 18:50:22 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
          Reply  Tue May 21 18:22:18 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
             Reply  Wed May 22 00:08:19 2013, Koji, Update, LSC, Sensing matrix scripts modified to include actuator calibration 
                Reply  Wed May 22 18:24:19 2013, Jenne, Update, SUS, Violin mode survey 
                Reply  Wed May 22 20:50:26 2013, Jenne, Update, LSC, Sensing matrix scripts don't calculate correctly 
                   Reply  Thu May 23 14:45:08 2013, Jenne, Update, LSC, Sensing matrix scripts calculations make more sense now SensMat_JenneMeas_23May2013.pngSensMat_KojiMeas_23May2013.png
                      Reply  Thu May 23 21:45:51 2013, Jenne, Update, LSC, Sensing matrix scripts now check for lockloss 
                Reply  Thu May 23 19:09:15 2013, Jenne, Update, LSC, Sensing matrix scripts modified to include actuator calibration SensMat_21May2013.png
          Reply  Wed May 22 18:07:36 2013, Jenne, Update, LSC, Kiwamu's sensing matrix measurement script revived 
             Reply  Thu May 23 00:49:13 2013, Jenne, Update, LSC, LSC filters loaded 
Message ID: 8632     Entry time: Thu May 23 19:09:15 2013     In reply to: 8611
Author: Jenne 
Type: Update 
Category: LSC 
Subject: Sensing matrix scripts modified to include actuator calibration 

After fixing up my calculations in my scripts, I have calculated the final PRMI sensing matrix (as measured very close to the violin frequency, so things may not be perfect). The data is from the file that Koji mentioned in his elog when he did the measurement:  elog 8611, sensematPRM_2013-05-22.12615.dat

Sensing Matrix, units = cts/meter, phase in degrees
 
            PRCL Mag   PRCL Phase    MICH Mag   MICH Phase  
AS55        1.064E+09   141.880      3.442E+09    11.929     
REFL11      3.474E+13  -108.372      4.316E+11   106.143     
REFL33      3.242E+11   -90.392      1.080E+10    81.037     
REFL55      3.373E+11   -92.060      1.394E+10    15.153 

SensMat_21May2013.png

In the plot, the little blobs on the ends of the 'sticks' are the error blobs.  Many of them are smaller than is really visible - this is good.  These errors come from measuring the lockin outputs several times while there is no drive to any optics, then the errors are propagated to each degree of freedom.  These errors do not incorporate any information about the precision of the actuator calibration, and they assume that the shape of all the sensor transfer functions are the same. 

If you look at the REFL11 and REFL33, it kind of seems like a miracle that we've ever been able to lock the full PRMI with the I&Q signals from either PD!

ELOG V3.1.3-